Suppose you have 30 CDs. You know that you have 11 more CDs than your friend. Write and solve an equation to find the number of CDs your friend has.
Answer by mananth(16949)

let your friend have x CDs
you have 11 more than your friend
so your friend has x+11 CD's
..
x+11 =30
x+11-11=30-11
x=19
Friend has 19 CDs